3. What is the time value of a call option costing $6 with a strike price of $32 and a spot price of $36?

4. What is the risk-free arbitrage profit available if a one-year option pair has a strike price of $60, a spot price of $62, a call premium of $5, a put premium of $2, and a risk-free rate of 4%?
Answer #1

1. =6-max(36-32,0)=2

2. S+P=2+62=64

C+xe^(-rt)=5+60*e^(-4%)=62.64737

S+P-C-Xe^(-rt)=64-62.64737=1.35263

Risk free arbitrage profit=1.35263